OpenStudy (anonymous):

Jesse paid $8.00 for a watermelon that weighed 16 pounds. What was the price per pound for the watermelon?

OpenStudy (a1234):

8/16 = x/1 8/16 = ?

OpenStudy (anonymous):

im in 7th grade.

OpenStudy (a1234):

Okay. You know how to do proportions and unit rates, right?

OpenStudy (lucaz):